OpenStudy (anonymous):

et f(x) = x + 7 and g(x) = x – 2. What is f(x) • g(x)? x^2 – 5x + 14 x^2 + 5x – 14 x^2 + 5x – 5 x^2 + 9x + 5

OpenStudy (anonymous):

f(x).g(x)=(x+7)(x-2)=x^2+7x-2x-14=x^2+5x-14 so b is right

jhonyy9 (jhonyy9):

so just you need multiplie these two functions